# ACEBench: Who Wins the Match Point in Tool Usage?

### v4

@@ Abstract
11footnotetext: Equal Contributions. Work was done during an internship at Huawei Noah’s Ark Lab. *Corresponding authors.33footnotetext: The code and datasets will be publicly available at GitHub.
@@ 1 Introduction
Large Language Models (LLMs), such as GPT-4 Achiam et al. (2023), have demonstrated exceptional performance across numerous natural language processing tasks Naveed et al. (2023); Qu et al. (2025); Mialon et al. (2023).
Studies have shown that incorporating tools can significantly expand LLM capabilities, particularly in specialized domains such as mathematics Das et al. (2024); Bulusu et al. (2024); Gou et al. (2023); Veerendranath et al. (2024), programming Xu et al. (2024), and reasoning Chen et al. (2022); Shao et al. (2022); Surís et al. (2023); Yang et al. (2023). On one hand, integrating tools into LLMs can enhance capabilities in multiple domains, for example, ToolTransformer Schick et al. (2023) enhances the ability of LLMs to solve complex problems by utilizing tools. On the other hand, adopting a tool usage paradigm can improve the robustness of the response and the transparency of the generation, thus increasing the explainability and trust of usersusers Schick et al. (2023), as well as improving the system’s adaptability. As this field continues to evolve, it is essential to comprehensively evaluate all aspects of tool usage, particularly in complex scenarios.
@@
While several studies have focused on evaluating tool usage Yan et al. (2024); Guo et al. (2024); Wang et al. (2024a); Qin et al. (2023); Wang et al. (2024b); Zhuang et al. (2023); Lu et al. (2024), there are still some shortcomings in the existing tool-use benchmarks. Firstly, existing benchmarks lack multi-turn dialogue evaluation in real-world scenarios. For example, the multi-turn dialogues in BFCL Yan et al. (2024) and HammerBench  Wang et al. (2024a) are composed of predefined fixed content combinations. Secondly, current tool-use benchmarks Qin et al. (2023); Guo et al. (2024); Huang et al. (2023); Li et al. (2023) lack fine-grained evaluation and personalized data assessment.
Additionally, existing benchmarks Qin et al. (2023); Guo et al. (2024); Wang et al. (2024b) ignore the assessment of special cases, or the evaluation methods are simplistic Yan et al. (2024), as user instructions in real life are not always perfectWang et al. (2024c). The model’s ability to recognize and handle these issues is also crucial for evaluation. Lastly, evaluation costs are high  Qin et al. (2023); Guo et al. (2024), as many studies rely on advanced large models for evaluation.
@@
Figure 1: Evaluation dataset construction pipeline: API Synthesis Module (left), Dialogue Generation Module (middle), Quality Inspection Module (right).
@@ 2 Related Works
Stable ToolBench Guo et al. (2024) addresses the issue of unstable external APIs by employing a virtual API server, but its dependence on large models for evaluation results in high costs and scalability challenges. BFCL Yan et al. (2024) introduces a benchmark for tool use in multi-turn dialogue scenarios. Yet, it assembles dialogues from fixed content, failing to capture the dynamic and adaptive nature of real-world interactions. Similarly, $\tau$-Bench Yao et al. (2024) evaluates language agents’ ability to engage with human users while adhering to domain-specific rules. Still, its narrow focus on just two scenarios limits its generalizability across diverse tasks. HammerBench Wang et al. (2024a) improves upon this by incorporating datasets derived from popular mobile applications and merging dialogues to simulate typical question-answer trajectories. However, like BFCL, its multi-turn dialogues are simplistic concatenations of pre-defined content, which do not reflect the complexities of real-world conversational dynamics.
In addition, some benchmarks Qin et al. (2023); Guo et al. (2024) rely on large language models (LLMs) for result evaluation, leading to high costs and unstable operations.
@@ 3.1 Dataset
We construct two versions of the dataset, one in Chinese and the other in English, ensuring an equal distribution of data types across both versions.
3.1.1 Data Generation
The data construction process is divided into two categories: (1) the construction of Agent data, as detailed in Appendix B.1, and (2) the construction of other types of data, which includes two primary steps: API synthesis and dialogue generation.
API Synthesis.
We use real APIs from various real-world scenarios as reference during synthesis to enhance authenticity. To ensure the stability of the data, we use synthetic APIs to construct the evaluation dataset, referencing real-world APIs as a guide. We employ a self-evolution approach by building a hierarchical API context tree to ensure the generated APIs cover a wide range of domains and functionalities Liu et al. (2024b). Initially, we extract relevant information from technical documents to guide the API generation. As the process progresses, the context tree is gradually expanded, ultimately ensuring the depth and breadth of the generated APIs. The left part in Figure 1 illustrates the generation of APIs.
Dialogue Construction. As shown in the middle part in Figure 1, we use two different dialogue generation pipelines built on the constructed API pool from which three to six candidate APIs are selected for each evaluation instance. For most cases, APIs are chosen randomly. However, for instances requiring specific functionality (e.g., similar APIs or multi-turn scenarios), advanced methods, including graph-based sampling Wang et al. (2024d), are used. Simple cases or those with predefined functionality use a template-based generation, where a single generator produces dialogues to ensure consistency. We employ a multi-agent dialogue pipeline for more complex scenarios, where three agents (user, assistant, and tool) role-play to simulate real-world interactions. Both pipelines are supported by carefully hand-crafted examples to ensure comprehensive coverage and diversity. A detailed description for special data construction is provided in Appendix B.2.
@@ 3.1.2 Multi-Stage Data Verification
To address issues like mismatched answers or ambiguous criteria, we have implemented a multi-stage verification process, shown on the right part of Figure 1.
@@
Human Quality Inspection. In the initial evaluation, the dataset remaining after automated quality inspection is assessed by three LLMs to assist human experts in data screening. Valid data is retained, while potentially problematic data is placed in the error candidate pool (approximately 20% of the dataset). These flagged entries undergo a two-step expert review process, where two experts independently assess and suggest modifications, and a third expert consolidates feedback, revising problem statements, API definitions, and answers. The revised data is re-evaluated and manually verified, and three rounds of optimization are performed to ensure a high-quality dataset.
@@ 3.1.3 Data Analysis
Figure 2: Distribution of APIs in terms of domains (Top 3 subcategories for each category).
@@
Number of turns and arguments. The test data in ACEBench covers a wide range of complexities. Specifically, we statistically analyzed the number of dialogue turns and the number of arguments in the called apis, which are visualized in Figure 4. The results show that the number of dialogue turns ranges from 1 to 8, encompassing most real-world scenarios.
These samples with varying numbers of turns and arguments further form a test suite that covers a broader range of difficulties, allowing ACEBench to conduct a more granular evaluation of different models.
Figure 5: Overview of evaluation process: The left represents ’Normal’ evaluation, which involves AST parsing for function and parameter checks. The middle illustrates ’Special’ evaluation, focused on error type matching. The right shows ’Agent’ evaluation, which assesses final state and intermediate results through user-model interaction.
@@ 3.2 Eval
In this section, we introduce the evaluation process, where different evaluation methods are applied based on the various data types. The inference prompt for evaluation can be found in Appendix D.
@@ 3.2.1 Normal Evaluation
As shown in the left part of Figure 5, we evaluate Normal Data by comparing the model’s function call outputs with the ground truth. We evaluate using AST parsing, starting by matching the function name to the target for exact correspondence, followed by verifying the consistency of parameter types, and finally confirming the accuracy of the parameter values. For cases with multiple possible correct answers, we use a candidate answer pool, and a match with any one of the candidates is considered correct. The Normal Data is evaluated using Accuracy, where a score of 1 indicates a complete match and 0 indicates a mismatch.
@@ 3.2.2 Special Evaluation
As shown in the middle part of Figure 5, the core goal of evaluating Special Data is to determine whether the model can accurately identify and point out the issue. Specifically, when the instruction is missing required parameters, the model needs to identify and indicate the missing parameter; when there is an error in the parameter, the model needs to correctly identify the error parameter; and when the task is irrelevant to the functions, the model needs to recognize that the function cannot process the request. In any of these cases, if the model correctly identifies and addresses the issue, the Accuracy is 1; otherwise, the Accuracy is 0.
@@ 3.2.3 Agent Evaluation
As shown in the right part of Figure 5, we evaluate the Agent Data by assessing the model’s ability to use tools in real-world interactions with humans, focusing on both result and process evaluation. There are two evaluation metrics:
@@ 4.1 Main results and analysis
Performance Variations Between Top Open-source and Closed-source Models. As shown in Figure 6, the performance of these top models is generally similar in basic ability tasks, with the GPT-4 series models leading in most tasks, demonstrating strong intelligence and adaptability. However, closed-source models outperform open-source models due to more refined internal optimizations, access to larger datasets, and potentially larger scales. A noticeable performance gap emerges in mid-to-high-level tasks, such as Special and Agent, as these tasks require higher-level abilities, such as handling abnormal situations, complex reasoning, strong multi-turn interaction and decision-making capabilities.
Figure 6: Performance of Top-3 closed/open-source models.
Loss of Generalization in Fine-Tuned Models. As shown in Table 3, models fine-tuned on specific datasets, such as Watt-Tool-8B Watt-AI (2024), xLAM-7B Liu et al. (2024c), and Hammer2.1-7BLin et al. (2024), exhibit a significant decline in performance on the Special dataset. This decline can primarily be attributed to the fact that while fine-tuning enhances a model’s performance on specialized tasks, it can also lead to a loss of generalization, making the model less effective on new or broader instruction-following tasks.
Table 3: The Accuracy of different models on Special Data (%).
Model Incomplete Error Irrelevant
Llama-3.1-8B-Instruct 29.0 20.0 14.0
Watt-Tool-8B 7.0 1.0 10.0
Qwen2.5-7B-Instruct 26.0 36.0 79.0
xLAM-7B-r 1.0 3.0 4.0
Llama-3.2-3B-Instruct 13.0 12.0 3.0
Hammer2.1-3B 0.0 3.0 0.0
@@
4.2 Error Analysis
Figure 7: Error type distribution across different model series on Normal Data.
Error Analysis of Special Data. As shown in Table 5, we identified two main types of model errors: The first type is "Error Detection", which refers to the model’s complete failure to detect issues in the user’s instructions or its inability to identify problems according to the prompt’s formatting requirements. The second type is "Error Correction," where the model detects the problem but provides unclear feedback. For example, the model might indicate that there is an issue, but fails to specify which parameter values are incorrect or what critical information is missing. Results show that most errors in special-type scenarios are caused by "Error Detection", highlighting a critical gap in the model’s problem-detection capabilities. This suggests that the model needs to learn not only simple tool invocation but also how to identify corresponding issues under imperfect instructions. Specific error examples for Special data can be found in Appendix G.2.
@@ 4.3 Further Analysis
Scaling Law.
We evaluated the performance of Qwen2.5-Coder (7B, 14B, 32B) and Qwen2.5-Instruct (3B, 7B, 14B, 32B, 72B) on the ACEBench dataset. As shown in Figure 8, the experimental results demonstrate that performance improves significantly across various tasks as the model size increases, with particularly strong results observed in high-complexity tasks. However, it is worth noting that as the model size continues to grow, the rate of performance improvement begins to slow down, especially between the 32B and 72B models. This indicates that while increasing the model parameters brings substantial performance gains initially, the marginal benefits of scaling up further decrease, making additional improvements more challenging.
Impact of Code Training. Based on the experimental results in Figure 8, we find that the score of Qwen2.5-Coder-7B is lower than the non-coder version, while the scores of Qwen2.5-Coder-14B and Qwen2.5-Coder-32B are higher than those of the non-coder versions of the same size. This indicates that code tuning becomes more impactful in improving model performance as the model size increases. For larger models (such as 14B and 32B), code tuning significantly enhances tool usage, especially in tasks that follow predefined rules and steps. However, for smaller models (such as 7B), despite code tuning, the differences between models are minimal, suggesting that smaller models have limited capacity and cannot fully leverage the advantages of code tuning.
@@ A.1 Data Categories Description
Agent Data refers to situations where completing a task in an environment built based on real-world scenarios typically requires multiple steps. In such cases, an advanced language model simulates the user role, replicating interactions found in real-world scenarios. This approach aims to evaluate the model’s ability to perform in complex interactions. We define a "multi-step" scenario as one in which the user appears only once in the entire conversation, while a "multi-turn" scenario is one where the user appears multiple times throughout the conversation.

### v5

@@ Abstract
11footnotetext: Equal Contributions. Work was done during Chen Chen’s internship at Huawei Noah’s Ark Lab22footnotetext: Corresponding authors.
@@ 1 Introduction
Large Language Models (LLMs), such as GPT-4 Achiam et al. (2023), have demonstrated exceptional performance across numerous natural language processing tasks Naveed et al. (2023); Qu et al. (2025); Mialon et al. (2023). Studies have shown that incorporating tools can significantly expand LLM capabilities, particularly in specialized domains such as mathematics Das et al. (2024); Bulusu et al. (2024); Gou et al. (2023); Veerendranath et al. (2024), programming Xu et al. (2024), and reasoning Chen et al. (2022); Shao et al. (2022); Surís et al. (2023); Yang et al. (2023). On one hand, integrating tools into LLMs can enhance capabilities in multiple domains, for example, ToolTransformer Schick et al. (2023) enhances the ability of LLMs to solve complex problems by utilizing tools. On the other hand, adopting a tool usage paradigm can improve the robustness of the response and the transparency of the generation, thus increasing the explainability and trust of usersusers Schick et al. (2023), as well as improving the system’s adaptability. As this field continues to evolve, it is essential to comprehensively evaluate all aspects of tool usage, particularly in complex scenarios.
@@
While several studies have focused on evaluating tool usage Yan et al. (2024); Guo et al. (2024); Wang et al. (2024a); Qin et al. (2023); Wang et al. (2024b); Zhuang et al. (2023); Lu et al. (2024), there are still some shortcomings in the existing tool-use benchmarks. Firstly, existing benchmarks lack multi-turn dialogue evaluation in real-world scenarios. For example, the multi-turn dialogues in BFCL Yan et al. (2024) and HammerBench  Wang et al. (2024a) are composed of predefined fixed content combinations. Secondly, current tool-use benchmarks Qin et al. (2023); Guo et al. (2024); Huang et al. (2023); Li et al. (2023) lack fine-grained evaluation and personalized data assessment. Additionally, existing benchmarks Qin et al. (2023); Guo et al. (2024); Wang et al. (2024b) ignore the assessment of special cases, or the evaluation methods are simplistic Yan et al. (2024), as user instructions in real life are not always perfectWang et al. (2024c). The model’s ability to recognize and handle these issues is also crucial for evaluation. Lastly, evaluation costs are high  Qin et al. (2023); Guo et al. (2024), as many studies rely on advanced large models for evaluation.
@@
Figure 1: Dataset construction pipeline. (a) Normal and Special data construction: API synthesis module (left), Dialogue generation module (right). (b) Agent Data Construction: include scenario construction, environment construction and question design.
@@ 2 Related Works
Stable ToolBench Guo et al. (2024) addresses the issue of unstable external APIs by employing a virtual API server, but its dependence on large models for evaluation results in high costs and scalability challenges. BFCL Yan et al. (2024) introduces a benchmark for tool use in multi-turn dialogue scenarios. Yet, it assembles dialogues from fixed content, failing to capture the dynamic and adaptive nature of real-world interactions. Similarly, $\tau$-Bench Yao et al. (2024) evaluates language agents’ ability to engage with human users while adhering to domain-specific rules. Still, its narrow focus on just two scenarios limits its generalizability across diverse tasks. HammerBench Wang et al. (2024a) improves upon this by incorporating datasets derived from popular mobile applications and merging dialogues to simulate typical question-answer trajectories. However, like BFCL, its multi-turn dialogues are simplistic concatenations of pre-defined content, which do not reflect the complexities of real-world conversational dynamics. In addition, some benchmarks Qin et al. (2023); Guo et al. (2024) rely on large language models (LLMs) for result evaluation, leading to high costs and unstable operations.
@@ 3.1 Dataset
We constructed two linguistically parallel versions of the dataset (Chinese and English), ensuring equal distribution of data types between them. The final dataset comprises 2,000 annotated entries.
3.1.1 Data Construction
The Normal and Special data are automatically generated by LLMs, whereas the Agent data is constructed by experts. Creation details for some data are provided in Appendix Section B.
Normal and Special Data Construction. We employ a fully automated LLM-based generation pipeline specifically designed for Normal and Special Data , as illustrated in Figure 1(a).
(1) API Synthesis. We use real APIs from various real-world scenarios as reference during construction to enhance authenticity. To ensure the stability of the data, we use synthetic APIs to construct the evaluation dataset, referencing real-world APIs as a guide. We employ a self-evolution approach by building a hierarchical API context tree to ensure the generated APIs cover a wide range of domains and functionalities Liu et al. (2024b). Initially, we extract relevant information from technical documents to guide the API generation. As the process progresses, the context tree is gradually expanded, ultimately ensuring the depth and breadth of the generated APIs.
(2) Dialogue Construction. We use two different dialogue generation pipelines built on the constructed API pool from which three to six candidate APIs are selected for each evaluation instance. For most cases, APIs are chosen randomly. However, for instances requiring specific functionality (e.g., similar APIs or multi-turn scenarios), advanced methods, including graph-based sampling Wang et al. (2024d), are used. Simple cases or those with predefined functionality use a template-based generation, where a single generator produces dialogues to ensure consistency. We employ a multi-agent dialogue pipeline for more complex scenarios, where three agents (user, assistant, and tool) to simulate real-world interactions. Both pipelines are supported by carefully hand-crafted examples to ensure comprehensive coverage and diversity.
Agent Data Construction. We implement a carefully curated human-expert construction framework specifically tailored for Agent Data generation, as shown in Figure 1(b).
(1) Scenario Construction. Through systematic abstraction of real-world interaction scenarios (such as food delivery services and telecommunication operations), we design functional modules with well-defined business semantics and specify each scenario’s core state variables (e.g., order status, account balance) and intrinsic property sets.
(2) Sandbox Environment Construction. We constructed an isolated sandbox environment with three core components: standardized functional interfaces with well-defined input/output specifications and preconditions, a dynamic attribute management system for real-time state transition monitoring, and an execution monitoring module that logs invocation processes.
(3) Question Design. Based on predefined multi-turn dialog specifications tailored to different scenarios, domain experts systematically crafted the conversational questions through an iterative annotation process.
Figure 2: Distribution of APIs in terms of domains (Top 2 subcategories for each category).
@@ 3.1.2 Multi-Stage Data Verification
To address issues like mismatched answers or ambiguous criteria, we have implemented a multi-stage verification process.
@@
Human Quality Inspection. In the initial evaluation, the dataset remaining after automated quality inspection is assessed by three LLMs to assist human experts in data screening. Valid data is retained, while potentially problematic data is placed in the error candidate pool. These flagged entries undergo a two-step expert review process, where two experts independently assess and suggest modifications, and a third expert consolidates feedback, revising problem statements, API definitions, and answers. The revised data is re-evaluated and manually verified, and three rounds of optimization are performed to ensure a high-quality dataset.
@@ 3.1.3 Data Analysis
Number of turns and arguments. The test data in ACEBench covers a wide range of complexities. Specifically, we statistically analyzed the number of dialogue turns and the number of arguments in the called apis, which are visualized in Figure 4. The results show that the number of dialogue turns ranges from 1 to 8, encompassing most real-world scenarios. These samples with varying numbers of turns and arguments further form a test suite that covers a broader range of difficulties.
@@
Figure 5: Overview of evaluation process: The left represents ’Normal’ evaluation: AST-based function and parameter verification. The middle illustrates ’Special’ evaluation: Imperfect instruction defect diagnosis. The right shows ’Agent’ evaluation: State transition analysis via user-model interaction
Table 3: The Accuracy of Models on Special Data
Model Incomplete Error Irrelevant
Llama-3.1-8B-Instruct 29.0 20.0 14.0
Watt-Tool-8B 7.0 1.0 10.0
Qwen2.5-7B-Instruct 26.0 36.0 79.0
xLAM-7B-r 1.0 3.0 4.0
Llama-3.2-3B-Instruct 13.0 12.0 3.0
Hammer2.1-3B 0.0 3.0 0.0
@@ 3.2 Eval
In this section, we introduce the evaluation process, and the inference prompt for evaluation can be found in Appendix D.
@@ 3.2.1 Normal Evaluation
As shown in the left part of Figure 5, we evaluate Normal Data by comparing the model’s function call output with the ground truth using AST parsing. For cases with multiple valid answers, we employ a candidate answer pool where matching any candidate constitutes correctness. Evaluation uses Accuracy metric (1=full match, 0=mismatch).
@@ 3.2.2 Special Evaluation
As illustrated in the middle section of Figure 5, the evaluation of Special Data primarily assesses the model’s capability in problem identification. Specifically, the model must: (1) detect and alert missing parameters, (2) accurately locate erroneous parameters, and (3) recognize task-function mismatches. For each case, Accuracy is scored as 1 if correctly identified, otherwise 0.
@@ 3.2.3 Agent Evaluation
As shown in the right part of Figure 5, we evaluate the agent’s capabilities by assessing the model’s proficiency in utilizing tools during human-agent interactions, employing gpt-4o as a user simulator for testing purposes. There are two evaluation metrics:
@@ 4.1 Main results and analysis
Loss of Generalization in Fine-Tuned Models. As shown in Figure 3, models fine-tuned on specific datasets, such as Watt-Tool-8B Watt-AI (2024), xLAM-7B Liu et al. (2024c), and Hammer2.1-7BLin et al. (2024), exhibit a significant decline in performance on the Special dataset. This decline can primarily be attributed to the fact that while fine-tuning enhances a model’s performance on specialized tasks, it can also lead to a loss of generalization, making the model less effective on new or broader instruction-following tasks.This phenomenon highlights the importance of balancing task-specific performance and generalization capability during model optimization.
@@
4.2 Error Analysis
Error Analysis of Special Data. As shown in Table 5, we identified two main types of model errors: The first type is "Error Detection", which refers to the model’s complete failure to detect issues in the user’s instructions or its inability to identify problems according to the prompt’s formatting requirements. The second type is "Error Correction," where the model detects the problem but provides unclear feedback. For example, the model might indicate that there is an issue, but fails to specify which parameter values are incorrect or what critical information is missing. Results show that most errors in special-type scenarios are caused by "Error Detection", highlighting a critical gap in the model’s problem-detection capabilities. This suggests that the model needs to learn not only simple tool invocation but also how to identify corresponding issues under imperfect instructions. Specific error examples can be found in Appendix G.2.
Figure 6: Error type distribution on Normal Data.
@@ 4.3 Further Analysis
Scaling Law. We evaluated the performance of Qwen2.5-Coder (3B, 7B, 14B, 32B) and Qwen2.5-Instruct (3B, 7B, 14B, 32B, 72B) on the ACEBench dataset. As shown in Figure 7, the experimental results demonstrate that performance improves significantly across various tasks as the model size increases, with particularly strong results observed in high-complexity tasks. However, it is worth noting that as the model size continues to grow, the rate of performance improvement begins to slow down, especially between the 32B and 72B models. This indicates that while increasing the model parameters brings substantial performance gains initially, the marginal benefits of scaling up further decrease, making additional improvements more challenging.
Table 6: Accuracy comparison of prompting strategies on English Normal Data (%).
Model Standard Condensed Minimal
Qwen2.5-3B-Instruct 34.5 31.8 27.8
Qwen2.5-7B-Instruct 48.5 47.5 45.5
Qwen2.5-14B-Instruct 56.3 54.0 47.5
Impact of Prompting Strategies. Prompt design significantly affects language model performance. We tested three strategies (see Appendix D.4):
(1)Standard Prompt: A comprehensive template designed to eliminate interference from information insufficiency, ensuring a fair evaluation.
(2)Condensed Prompt: A compact version retaining core instructions, testing performance with reduced but sufficient guidance.
(3)Minimal Prompt: A highly abbreviated form (e.g keywords) to assess the model’s ability to infer tasks from ultra-concise input.
The experimental results in Table 4 demonstrate that models utilizing standard prompt templates achieve the highest overall accuracy. This optimal performance can be attributed to the rigorous formatting specifications in standard prompts, which effectively mitigate interference from extraneous variables. These empirical findings establish a positive correlation between prompt standardization and model performance, providing key insights for future prompt engineering: enhancing the standardization of function-calling prompts with explicit formatting requirements can significantly improve execution accuracy.
@@ A.1 Data Categories Description
Agent Data refers to scenarios where completing a task in an environment modeled after real-world situations typically requires multi-step collaboration. In this study, we employ the GPT-4o language model to simulate user roles and replicate real-world interaction processes, thereby evaluating the model’s performance in complex interactive settings. The key scenarios are defined as follows:
Multi-step Scenario: The user participates in only a single interaction throughout the entire dialogue flow.
Multi-turn Scenario: The user engages in multiple interactions across the entire dialogue cycle.
Agent Data currently encompasses the following fundamental scenarios:
(1)Mobile Application Simulation. The mobile application scenario provides digital lifestyle functionalities including communication services, integrated reminder and memo management systems, and alarm configuration capabilities. This environment simulates core smartphone operations with particular attention to notification handling and scheduling precision.
(2)Food Delivery Platform. This scenario simulates the core functionalities of a food delivery platform, primarily including merchant search, product browsing and ordering, order status tracking, and cancellation processing. The system implements essential operational procedures from merchant selection to order completion, supporting users throughout the entire food ordering experience.
(3)Financial Services Scenario. This module provides fundamental banking service simulations, primarily including: deposit/withdrawal transactions, account balance inquiries, fund transfers, and other routine banking operations, while also supporting loan applications and repayment processes. The system maintains detailed transaction records and can generate basic financial statements, replicating the core services of real banking systems.
(4)Travel Booking Platform. This scenario simulates a standard ticketing system, enabling users to complete end-to-end operations including flight/train ticket inquiry, booking, payment, rescheduling, and cancellation. The system incorporates fare checking, seat selection, and order management functionalities, capable of handling itinerary changes and related ticket adjustments. It covers the complete user journey from search to ticket issuance.
Our team is actively working on designing more functional scenarios to enhance the platformin Agent Data.